# Vector interaction bounds in NJL-like models from LQCD estimated curvature of the chiral crossover line

### v1

@@ III.3 $\mu_{S}$ variation of $\kappa_{4}^{B}$
Figure 4: The curvature coefficient for different values of $\mu_{S}$ in Model II with the same parameter sets used in the above graph.
Although, in scenarios considered in LQCD, the curvature coefficient $\kappa_{4}^{B}$ are obtained to be zero, we have obtained a significant dependence on $\mu_{S}$. Fig. 4 represents the models’ predictions on $\kappa_{4}^{B}$ as a function of $\mu_{S}$. The overall behaviour of $\kappa_{4}^{B}$ as a function of $\mu_{S}$ is the same (increasing) for both Model-I and Model-II with the corresponding attractive interactions $(G_{V}^{-}/g_{V}^{-})$ producing a higher curvature as compared to repulsive ones for a given value of $\mu_{S}$. This observation might provide us with a better understanding of the effect of $\mu_{S}$ as well as the vector interaction on the crossover line once explored in LQCD.
@@ III.4 Effect of vector interaction on the critical endpoint
Figure 5: Effect of the vector interactions on the QCD phase line for $\mu_{S}=0$. Blue and magenta lines denote the estimation from the upper and lower bounds of the flavor-dependent vector interaction $g_{V}$, respectively. Soft blue and soft orange denote the same for flavor-independent scenario, i.e., with nonzero $G_{V}$. The red line denotes the limiting case with no vector interaction, $g_{V}=G_{V}=0$. The circles denote the CEP for the corresponding lines. Closed and open symbols are for Model-I and Model-II, respectively. (The line overlap for $G^{+}_{V}$ and $g^{+}_{V}$ at low $\mu_{B}$).
@@
In Fig. 5, we show the evaluated phase lines for constrained values of $G_{V}$ and $g_{V}$, with the corresponding CEPs indicated by circle symbols. The exact locations of the CEPs are provided in Table 4. As the repulsive vector interaction flattens the phase boundary, the CEP is expected to shift rightward for the positive upper limits of both $G_{V}$ and $g_{V}$, resulting in a corresponding decrease in $T^{\text{CEP}}$, as seen for the $g_{V}^{+}$ line. It is important to note that in Model-I, all three quark flavors contribute to the shifted chemical potential $\tilde{\mu}$. In contrast, Model-II includes flavor-dependent contributions, where only individual quark flavors affect $\tilde{\mu}$. As a result, the phase boundary is more compressed in the flavor-independent scenario. These results align with those in Ref. [19]. For the negative lower values of $G_{V}$ and $g_{V}$, the CEP shifts leftward, with an associated increase in $T^{\text{CEP}}$. It should be emphasized that the locations of these CEPs do not contradict predictions from Lattice QCD [7] or FRG-based studies [8], but instead offer a baseline within this effective model framework. Future constraints on the CEP location from first-principle QCD calculations may help narrow or exclude the range of coupling strengths explored here.
@@ IV Summary and Outlook
In this work, we investigate the effect of vector-type interactions within the $2+1$-flavor Nambu–Jona-Lasinio (NJL) model, considering both flavor-independent and flavor-dependent interactions with coefficients $G_{V}$ and $g_{V}$, respectively. These interactions arise only in the finite-density regime, making it difficult to fix the strengths of $G_{V}$ and $g_{V}$ from the QCD vacuum properties. Previous attempts to constrain these coefficients using LQCD-estimated curvature coefficients of the crossover line were limited due to large uncertainties in the lattice results. Here, we revisit this issue and utilize the precise lattice results for $\kappa_{2}^{B}$ for the cases $\mu_{S}=0$ and $n_{s}=0$ to constrain $G_{V}$ and $g_{V}$. We varied these coefficients in units of $G_{S}$ over a range of negative to positive values and evaluated $\kappa_{2,4}^{B}$ for the chiral crossover line. Negative values of $G_{V}$ and $g_{V}$ produce larger values of $\kappa_{2}^{B}$, which decrease as $G_{V}/G_{S}$ and $g_{V}/G_{S}$ increase towards positive values and further. This approach provides a tighter constraint on the range of these interaction coefficients. In the $\mu_{S}=0$ case, the allowed ranges of interaction strength are $\left[-0.117,0.212\right]$ and $\left[-0.296,0.552\right]$ (in units of $G_{S}$) for the flavor-independent and flavor-dependent interactions, respectively. Conversely, the $n_{s}=0$ case provides narrower ranges: $\left[-0.038,0.334\right]$ for $G_{V}/G_{S}$ and $\left[-0.076,0.667\right]$ for $g_{V}/G_{S}$. As the estimates for $\kappa_{4}^{B}$ are consistent with lattice results and remain within variances [4] for a wide range of $G_{V},g_{V}$, they do not further constrain the vector interaction strength. At this juncture, we want to iterate that the bound on $G_{V},g_{V}$ we get for the scenarios $\mu_{S}=0$ and $n_{s}=0$ are different, which necessitates the consideration of a more realistic interaction strength which depends on $\mu$.
@@
We find that the improved lattice data on $\kappa_{2}^{B}$ provides narrower bounds on the vector interactions and can accommodate both attractive and repulsive vector interaction within the model framework. We test the model’s prediction in different physical scenarios, such as zero strangeness or the strangeness neutrality condition. $\kappa_{2}^{B}$ is always smaller in the latter case for both increasing $G_{V}$ and $g_{V}$, except at higher values of $G_{V}$ (flavor-independent scenario), where we observe a crossing. However, the bound is always wider on $g_{V}$ as compared to $G_{V}$, which can be explained within the model framework. The models’ predictions on the fourth-order curvature coefficient $(k_{4}^{B})$ match with the corresponding LQCD results, which can be termed as consistent with zero with the error bars.
@@
In the NJL model, the mixing between light and strange quarks primarily arises from the ’t Hooft determinant term, introduced to take into account the QCD axial anomaly. The flavor-independent vector interaction also couples the light and strange quark sectors through their net densities. In this work, we have proposed a novel method to separate out the contribution of the flavor-independent vector interaction to flavor mixing by investigating the effect of the strangeness chemical potential, $\mu_{S}$, on $\kappa_{2}^{B}$, which was first proposed in Ref. [27]. These flavor mixings between light and strange sectors are responsible for the decrease in $\kappa_{2}^{B}$ as $\mu_{S}$ increases. Our findings are in general agreement with those from Ref. [27] We also observe that at a specific value of $\mu_{S}$, the effect of the vector interactions become negligible for both Model-I and Model-II and match with those for $G_{V}=g_{V}=0$ limit. This value of $\mu_{S}$ is larger for Model-II, signifying weak flavor mixing in the flavor-dependent coupling as compared to the flavor-independent one. Such interesting characteristics provide us with a better understanding of the types of vector interactions and possible signatures to distinguish them.
@@
To assist the previous understanding, we have studied the case where the ’t Hooft determinant term is zero. This has enabled us to explicitly demonstrate the impact of vector interactions on the variation of $\kappa_{2}^{B}$ and allow for direct assessment of the flavor mixing driven by this flavor-independent vector interaction. In this limit, there is no variation of $\kappa_{2}^{B}$ in model II, indicating no mixing effect for the flavor-dependent vector interaction.
@@
For the final part of our study, we find the location of the CEP for the allowed range of vector interactions. The location of CEP is heavily impacted by the strength of the vector interaction. We find that the CEP exists for the negative lower bounds of the vector interactions and disappears for the positive upper bounds in both the Model-I and Model-II. Although the model’s predictions on the location of CEP are consistent with the existing LQCD bound [7, 8], one should keep in mind that such observations can, in principle, depend on multiple factors, capturing all of which remains a task beyond such simple effective model calculations.
@@ Acknowledgment
D.B. is supported by the Department of Science and Technology, Government of INDIA under the SERB National Post-Doctoral Fellowship Reference No. PDF/2023/001762. C.A.I. acknowledges Goethe University’s R3 Career Support, supported by Johanna Quandt Young Academy @@@@Goethe (JQYA) and the support by the State of Hesse within the Research Cluster ELEMENTS with project ID 500/10.006. He also thanks ECT* and CRC for support at the Workshop “Spin and quantum features of QCD plasma” during which this work has been completed.
@@ Appendix A Variation of $\kappa_{4}^{B}$ with $G_{V}$ and $g_{V}$
Figure 6: The curvature coefficients, $\kappa_{4}^{B}$ as a function of $G_{V}/G_{S}$ (upper panel) $g_{V}/G_{S}$ (lower panel), respectively. The LQCD estimation of the same is $0.001(7)$ [4].
@@ Appendix B $\kappa_{2}^{S}$ as a function of $G_{V}$ and $g_{V}$
Figure 7: The curvature coefficients, $\kappa_{2}^{S}$ as a function of $G_{V}/G_{S}$ (upper panel) $g_{V}/G_{S}$ (lower panel), respectively. The LQCD estimation of the same is $0.017(5)$ [4].

### v2

@@ III.3 $\mu_{S}$ variation of $\kappa_{4}^{B}$
Figure 4: The curvature coefficient for different values of $\mu_{S}$ in Model-II with the same parameter sets used in the above graph.
Although, in scenarios considered in LQCD, the curvature coefficient $\kappa_{4}^{B}$ are obtained to be zero, we have obtained its significant dependence on $\mu_{S}$. Fig. 4 represents the models’ predictions on $\kappa_{4}^{B}$ as a function of $\mu_{S}$. The overall behaviour of $\kappa_{4}^{B}$ as a function of $\mu_{S}$ is the same (increasing) for both Model-I and Model-II with the corresponding attractive interactions $(G_{V}^{-}/g_{V}^{-})$ producing a higher curvature as compared to repulsive ones for a given value of $\mu_{S}$. This observation might provide us with a better understanding of the effect of $\mu_{S}$ as well as the vector interaction on the crossover line once explored in LQCD.
@@ III.4 Effect of vector interaction on the critical endpoint
Figure 5: Effect of the vector interactions on the QCD phase line for $\mu_{S}=0$. Blue and magenta lines denote the estimation from the upper and lower bounds of the flavor-dependent vector interaction $g_{V}$, respectively. Soft blue and soft orange denote the same for flavor-independent scenario, i.e., with nonzero $G_{V}$. The red line denotes the limiting case with no vector interaction, $g_{V}=G_{V}=0$.
@@
In Fig. 5, we show the evaluated phase lines for constrained values of $G_{V}$ and $g_{V}$, with the corresponding CEPs indicated by circle symbols. The exact locations of the CEPs are provided in Table 4. As the repulsive vector interaction flattens the phase boundary, the CEP is expected to shift rightward for the positive upper limits of both $G_{V}$ and $g_{V}$, resulting in a corresponding decrease in $T^{\text{CEP}}$, as seen for the $g_{V}^{+}$ line. It is important to note that in Model-I, all three quark flavors contribute to the shifted chemical potential $\tilde{\mu}$. In contrast, Model-II includes flavor-dependent contributions, where only individual quark flavors affect $\tilde{\mu}$. As a result, the phase boundary is more compressed in the flavor-independent scenario. These results align with those in Ref. [21]. For the negative lower values of $G_{V}$ and $g_{V}$, the CEP shifts leftward, with an associated increase in $T^{\text{CEP}}$. It should be emphasized that though the locations of these CEPs are different from those found using functional methods [36, 37], none of them contradicts predictions from Lattice QCD [7, 8], and they offer a baseline within this effective model framework. However, another important ingredient of the model, the strength of the ’t Hooft determinant term $(K)$, plays crucial role on determining the location of the CEP [21]. At present, there are some arbitrariness for the value of $K$ (Table 1). A precise determination of $K$, possibly by following the technique mentioned in Ref. [35], along with future constraints on the CEP location from first-principle QCD calculations, may help narrow down the range of the vector interaction strengths explored here.
@@
We want to mention that the obtained locations of the CEP with the other parameter set (set II) are consistent with earlier works  [63, 64, 65]. The effect of vector interactions on the CEP locations is qualitatively similar to that of set I. As $T^{\rm CEP}$ from set II is significantly larger compared to the one from parameter set I, this necessitates a stronger repulsive interaction to make the CEP disappear from the phase diagram. As the allowed range of $G_{V}$ in our work is almost similar for both parameter sets, the CEP does not disappear from the phase diagram for set II even for the repulsive interaction. It should be emphasized that the predictions of CEP locations within the NJL-like model are very sensitive to the model parameters [66]. For example, stronger values of the eight-quark coupling $g_{1}$ would push the temperatures associated with CEP to even larger values [67]. In the present study, we do not intend to quantify the absolute location of the CEPs; rather, we focus on their shifts with varying vector interaction strength.
@@ IV Summary and Outlook
In this work, we investigate the effect of vector-type interactions within the $2+1$-flavor Nambu–Jona-Lasinio (NJL) model, considering both flavor-independent and flavor-dependent interactions with coupling strengths $G_{V}$ and $g_{V}$, respectively. In general, the strengths of these coupling constants can be fixed from the masses of the vector mesons [19, 12]. However, in an effective model scenario, such an exercise is less robust than the determination of the scalar coupling constant. On the other hand, in a dense medium, vector interactions can be induced due to their direct coupling with the number density operator [21, 22]. Such induced vector interactions arise only in a finite-density regime, making it difficult to fix the strengths of induced $G_{V}$ and $g_{V}$ from the QCD vacuum properties. Previous attempts to constrain these coupling strengths using LQCD-estimated curvature coefficients of the crossover line were limited due to large uncertainties in the lattice results.
@@
Here, we revisit this issue and utilize the precise lattice results for $\kappa_{2}^{B}$ for the cases $\mu_{S}=0$ and $n_{s}=0$ to constrain both $G_{V}$ and $g_{V}$. We varied these couplings in units of $G_{S}$ over a range of negative to positive values and evaluated $\kappa_{2,4}^{B}$ for the chiral crossover line. Negative values of $G_{V}$ and $g_{V}$ produce larger values of $\kappa_{2}^{B}$, which decrease as $G_{V}/G_{S}$ and $g_{V}/G_{S}$ increase towards positive values and further. This approach provides a tighter constraint on the range of these interaction strengths. In the $\mu_{S}=0$ case, the allowed ranges of interaction strength are $\left[-0.117,0.212\right]$ and $\left[-0.296,0.552\right]$ (in units of $G_{S}$) for the flavor-independent and flavor-dependent interactions, respectively. Conversely, the $n_{s}=0$ case provides narrower ranges: $\left[-0.038,0.334\right]$ for $G_{V}/G_{S}$ and $\left[-0.076,0.667\right]$ for $g_{V}/G_{S}$. As the estimates for $\kappa_{4}^{B}$ are consistent with lattice results and remain within variances for a wide range of $G_{V},g_{V}$, they do not further constrain the vector interaction strength. At this point, we want to iterate that the bound on $G_{V},g_{V}$ we get for the scenarios $\mu_{S}=0$ and $n_{s}=0$ are different, which necessitates the consideration of a more realistic interaction strength which depends on $\mu$.
We find that the improved lattice data on $\kappa_{2}^{B}$ provides bounds on the vector interactions accommodating both attractive and repulsive vector interaction within the model framework. Such findings highlight the fact that lattice-calculated bounds on curvature coefficients, which has been previously used to constrain induced vector interactions [24, 25, 26, 27, 28], allow for positive and negative induced vector interactions. However, such a bound should be tested against other phenomenological consequences of the vector interaction; for instance, neutron star observations favor a repulsive (positive) vector interaction, which is required to generate an equation of state stiff enough to support two-solar-mass stars [68, 69]. On the other hand, attractive vector interaction can lead to larger values of the susceptibilities and thus mechanical instability at high density [29], particularly with a constant vector interaction. Such uncertainties make the strength and sign of the vector interaction a critical issue. Future lattice QCD determinations of susceptibilities at larger baryon densities will therefore be invaluable in constraining the magnitude of ${G_{V}}~\text{and}~{g_{V}}$.
It should also be mentioned that the masses of vector mesons allow only repulsive interactions [19]. However, one needs to take into account both the quark loop and the pseudoscalar meson loops in the calculation of vector meson self-energy [19]; otherwise, the vector interaction may appear with the wrong sign [70]. Using vector meson masses, Ref.[19] extracted a value of $G_{V}/G_{S}\sim 1.5$, whereas Ref.[22] predicts the ratio to be $\sim 0.6$, thus allowing for a wide range depending on the model parameters. The upper limits of the bounds on the induced vector interaction that we find are therefore not unreasonable with such a wide range, if one attempts to determine the vector meson masses from an induced vector interaction.
@@
While testing the model’s predictions under the zero strangeness or strangeness neutrality condition, we observe that $\kappa_{2}^{B}$ is always smaller in the latter case for both increasing $G_{V}$ and $g_{V}$, except at higher values of $G_{V}$ (flavor-independent scenario), where we observe a crossing. However, the bound is always wider on $g_{V}$ as compared to $G_{V}$, which can be explained within the model framework. The models’ predictions on the fourth-order curvature coefficient $(k_{4}^{B})$ match with the corresponding LQCD results, which can be termed as consistent with zero with the error bars. We intentionally avoided the Polyakov loop dynamics to focus solely on the chiral dynamics. However, a similar analysis has been performed in the presence of Polyakov loop dynamics in Refs. [24, 25, 26, 27], where the curvature coefficients $\kappa_{2}^{B}$ were found to be very close to our NJL estimate. Since the curvature is calculated for the transition line up to $\mu_{B}/T\leq 1$, we expect the allowed range of the vector interaction strength not to change significantly in the presence of a background gauge field. Nevertheless, this is certainly an interesting aspect to explore and verify explicitly, which we plan to take up in future work.
@@
In the NJL model, the mixing between light and strange quarks primarily arises from the ’t Hooft determinant term, introduced to take into account the QCD axial anomaly. The flavor-independent vector interaction also couples the light and strange quark sectors through their net densities, thereby contributing to flavor mixing. In this work, we have proposed a novel method to separate out the contribution of the flavor-independent vector interaction to flavor mixing by investigating the effect of the strangeness chemical potential, $\mu_{S}$, on $\kappa_{2}^{B}$. These flavor mixings between light and strange sectors are responsible for the decrease in $\kappa_{2}^{B}$ as $\mu_{S}$ increases. Our findings are in general agreement with those from Ref. [35]. We also observe that at a specific value of $\mu_{S}$, the effect of the vector interactions becomes negligible for both Model-I and Model-II and match with those for $G_{V}=g_{V}=0$ limit. This value of $\mu_{S}$ is larger for Model-II, signifying weak flavor mixing in the flavor-dependent coupling as compared to the flavor-independent one. Such interesting characteristics provide us with a better understanding of the types of vector interactions and possible signatures to distinguish them.
@@
To elaborate on the contribution from different mixing effects, we have studied the case where the ’t Hooft determinant term is taken to be zero. This has enabled us to explicitly demonstrate the impact of vector interactions on the variation of $\kappa_{2}^{B}$ and allows for a direct assessment of the flavor mixing driven by this flavor-independent vector interaction. In this limit, there is no variation of $\kappa_{2}^{B}$ in model II, indicating no mixing effect for the flavor-dependent vector interaction.
@@
For the final part of our study, we find the location of the CEP for the allowed ranges of vector interactions. The location of CEP is heavily impacted by the strength of the vector interaction. We find that the CEP exists for the negative lower bounds of the vector interactions and disappears for the positive upper bounds in both the Model-I and Model-II. Although the model’s predictions on the location of CEP are consistent with the existing LQCD bound, one should keep in mind that such observations can, in principle, depend on multiple factors, capturing all of which remains a task beyond such simple effective model calculations.
@@ Acknowledgment
D.B. is supported by the Department of Science and Technology, Government of INDIA under the SERB National Post-Doctoral Fellowship Reference No. PDF/2023/001762. C.A.I. acknowledges Goethe University’s R3 Career Support, supported by Johanna Quandt Young Academy @@Goethe (JQYA) and the support by the State of Hesse within the Research Cluster ELEMENTS with project ID 500/10.006. He also thanks ECT* and CRC for support at the Workshop “Spin and quantum features of QCD plasma” during which this work has been completed.
@@ Appendix A Variation of $\kappa_{4}^{B}$ with $G_{V}$ and $g_{V}$
In Fig. 6, we present the $4$th order curvature coefficient, $\kappa_{4}^{B}$, as a function of $G_{V}$ (Model-I) and $g_{V}$ (Model-II), respectively. We explored it for the same ranges of $G_{V}$ and $g_{V}$ that we used to explore both $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$. It remains zero for a repulsive interaction (positive $G_{V}/g_{V}$) but starts to decrease as the strength of the attractive interaction (negative $G_{V}/g_{V}$) increases. The LQCD estimates of $4$th order curvature coefficient for $\mu_{S}=0$ and $n_{s}=0$ are $0.001(7)$ and $0.004(6)$, respectively [4]. Due to the large uncertainties in LQCD estimates, we cannot further constrain $G_{V}/g_{V}$ beyond the limits obtained from $\kappa_{2}^{B}$s.
Figure 6: The curvature coefficients, $\kappa_{4}^{B}$ as a function of $G_{V}/G_{S}$ (upper panel) $g_{V}/G_{S}$ (lower panel), respectively. The LQCD estimations of the same are $0.001(7)$ and $0.004(6)$ for $\mu_{S}=0$ and $n_{s}=0$, respectively [4].
@@ Appendix B $\kappa_{2}^{S}$ as a function of $G_{V}$ and $g_{V}$
The curvature coefficient $\kappa_{2}^{S,\mu_{B}=0}$ as a function of both $G_{V}$ and $g_{V}$ are presented in Fig. 7. The effect of $G_{V}/g_{V}$ on $\kappa_{2}^{S,\mu_{B}=0}$ is comparatively weaker than on $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$. The dominant effect of both vector interactions $G_{V},g_{V}$ on the light quark chiral transitions propagates through the respective number densities. For $\mu_{B}=0$, the light quark number densities get minor modifications due to $n_{s}\neq 0$ leading to $\tilde{\mu}_{B}\neq 0$ in Model-I; whereas they remain zero in Model-II. Hence, in Model-II, $g_{V}$ modifies only the strangeness chemical potential, which influences the chiral transition through the ’t Hooft determinant term, resulting in a monotonically decreasing $\kappa_{2}^{S,\mu_{B}=0}$ as a function of $g_{V}$. On the other hand, Model-I is a little more subtle, as it induces a finite $\mu_{B}$, leading to a decrease in the crossover temperature. This reduces the effect of repulsive interaction (positive $G_{V}$) resulting in a slow variation of $\kappa_{2}^{S,\mu_{B}=0}$ as a function of $G_{V}$. On the other hand, for attractive interaction (negative $G_{V}$), the curvature increases at a faster rate. From the upper panel of Fig. 7, the difference in slope in the positive and negative range of $G_{V}$ is clearly visible, where we have used identical $G_{V}/g_{V}$ ranges as used in Fig. 1.
To compare Model-I and Model-II, we present the curvature $\kappa_{2}^{S,\mu_{B}=0}$ as a function of $G_{V}/g_{V}$ within a given range for both models in Fig. 8. As already described, in Model-II, it has a monotonic decreasing behavior, whereas the effect of $G_{V}$ gets suppressed in the positive range and enhanced in the negative range for Model-I. It is to be noted that $\mu_{B}=0$ leads to $G_{V}\simeq g_{V}$, which is evident from Eq. (7).
From Fig 7, one might obtain an upper bound of $g_{V}$ in Model-II, but as presented in Table 2, one can consider a different parameter set, mainly a different $K$ and shift these curves upward resulting in a different bound. Hence, we conclude that without a proper constraint on $K$, which can be obtained following the procedure presented in Ref. [35], one cannot obtain a meaningful bound on the vector interactions from $\kappa_{2}^{S,\mu_{B}=0}$.
Figure 7: The curvature coefficients, $\kappa_{2}^{S,\mu_{B}=0}$ as a function of $G_{V}/G_{S}$ (upper panel) $g_{V}/G_{S}$ (lower panel), respectively. The LQCD estimation of the same is $0.017(5)$ [4].
Figure 8: The curvature coefficients, $\kappa_{2}^{S,\mu_{B}=0}$ as a function of the vector interactions for both models.